# Advancing Brainwave-Based Biometrics: A Large-Scale, Multi-Session Evaluation

### v1

@@ 4.2.1. Effect of the Number of Test Subjects on EER:
When comparing the results of different experiments using metrics such as EER, number of subjects in the test is typically reported as metadata. A larger test set generates more impostor scores in zero-effort scenarios, making it valuable to investigate the relationship between the number of subjects in the test set and the resulting error. This analysis provides insights for better comparisons of results across different studies or our follow-up experiments with smaller test sets, such as when evaluating performance using headset-specific recording. To achieve this, we randomly select $N$ subjects, where $N$ ranges from 2 to 99 (starting at $N=2$ since at least one genuine user and one impostor are required). For each value of $N$, we compute the EER, repeat the procedure 50 times, and report the average EER per $N$.
@@
Figure 2 presents the results of the experiment. The analysis revealed no significant relationship between the number of subjects and the EER. These findings align with observations in face recognition systems. For example, Friedman et al. (Friedman et al., 2022) investigated how test size affects identification and authentication tasks. They reported a linear decline in identification performance with the logarithm of enrollment size but found that the EER for authentication tasks remained stable. This stability occurs because the EER is based on the Receiver Operating Characteristic (ROC) curve, which balances the FAR and FRR derived from impostor and genuine score distributions. While larger subject pools better define these distributions, they do not significantly influence the ROC. Therefore, we can conclude that having a higher number of subjects in the test set does not necessarily make authentication more challenging. However, both scenarios indicate that test sets with a small number of participants exhibit high standard deviations in EER, which can lead to unreliable or misleading results. This underscores the importance of using larger test sets to obtain more dependable evaluations.
Figure 4. t-SNE Visualization of user identity embeddings, where a color is assigned to each of the 100 subjects in the test set. Due to the large number of subjects, some colors may repeat.
4.2.2. Correlation Between EER and Time Interval:
@@
To investigate the permanence, we first visualize the distribution of test session intervals in Figure 3. The test data includes a total of 1,699 unique sessions available for 100 test subjects, where the first session (or the oldest session) per subject is considered as the enrollment session (100 sessions). The results show that the highest concentration of sessions occurred within the first week, with 129 sessions (8%) recorded. By the end of the first month (weeks 0 to 3), this number increased to 560 sessions (35%), indicating that one-third of all sessions were conducted in the first month. The second month (weeks 4 to 7) accounted for 398 sessions (24.8%), while the third month (weeks 8 to 11) added 201 sessions (12.5%). Beyond the third month, session frequency declined further, with 182 sessions (11.3%) recorded from Month 4 to the end of the first year. In the second year, session numbers were 89 sessions (5.5%), and remained similar at 71 sessions (4.4%) in the third year. The remaining years combined accounted for 4.8%.
@@
To investigate the correlation between time intervals between sessions and authentication error, we extracted all possible pairs of sessions within specific time intervals and calculated the EER for each. Figure 3 shows the results, demonstrating a positive correlation between time interval and error rate, as expected. Verifying identity becomes harder over time. The EER after one day is 7.7%, increasing to 8.6% after seven days. However, within the second month, the EER rises to 10.4%, which remains lower than the previously reported average of 10.78%, indicating stable performance over a reasonable time interval. Beyond the second month, the EER gradually increases, reaching 19.69% after one year. Performance remains relatively stable in the second and third years. However, in the fourth and fifth years, the performance deteriorates significantly, approaching random baseline levels with EER values of 40% and 75%, respectively. Therefore, it is necessary to update the enrollment set at shorter time intervals. This can be achieved by using verification data from successful login attempts to update the enrollment set, ensuring that the process does not impose additional overhead on the user.
@@
Figure 5. t-SNE visualization of embeddings, where each color represents a different EEG recording headset used for data collection.
4.2.3. Visualize Embedding Space:
@@
To better understand how well users are represented in distinct clusters, we transform the 128-dimensional embeddings to a 2D space using t-Distributed Stochastic Neighbor Embedding (t-SNE) and visualize the results. While some noise is present, the Figure 4 shows that the metric learning approach successfully separates subjects into distinguishable clusters, demonstrating the model’s ability to capture individual differences. However, we observe two separate colonies within the embedding space, leading us to hypothesize that these colonies could be due to differences in the hardware used during data capture. This hypothesis is confirmed by Figure 5, which reveals that the data captured using BioSemi hardware is located separately from the rest, further illustrating the significant influence of hardware on the embedding distribution. The effect of hardware on the results will be discussed in the next section.
@@
Figure 6. t-SNE visualization of embeddings for 20 randomly selected subjects, with each subject assigned a unique color for a clear distinction. Enrollment samples are depicted as circles without borders, while verification samples are represented by various shapes with black borders. Also, each unique shape corresponds to a distinct session for verification samples.
@@
Figure 7. Magnified view of the right-middle region of Figure 6, highlighting the green subject. Each session for the highlighted subject is assigned a unique color, while all other subjects are shown in gray for context.
@@
To investigate the embedding space while accounting for color and spatial limitations, we randomly selected 20 subjects to prevent overcrowding the embedding space with excessive data. [^7] BioSemi subjects were excluded to avoid large gaps in the plot and enhance detail visibility. Figure 6 illustrates the embedding space for these selected subjects, demonstrating that the feature extraction method effectively identifies subject identities. However, a session effect is evident, where samples from the same session tend to form distinct clusters. To better visualize this phenomenon, the right-middle region of Figure 6 is magnified in Figure 7, with separate colors indicating different sessions, further confirming the session-based clustering. The session effect on clustering could primarily arise from environmental noise, changes in hair length, slight variations in the positioning of brainwave channels on the head, hardware differences, and natural changes in brainwave patterns over time due to factors such as aging, personal experiences, or mental states (e.g., stress or focus levels) throughout sessions.
4.2.4. Impact of Different Headsets on Performance
@@
Due to differences in recording hardware technologies and channel positions in EEG recording devices, further investigation is required to understand their effect on brainwave authentication. Specifically, we aim to address three questions: (1) To what extent can a model trained on one hardware be used for the authentication of user data recorded by another hardware? (2) Can models trained on multiple hardware types outperform headset-specific models? (3) What is the performance of cross-device authentication when enrollment and verification data come from different hardware?
@@
Cross Model: To address the first question, we trained models using data from each specific headset and grouped the test data based on hardware to evaluate performance across models. The results, summarized in Table 2, show that headset-specific models perform well only when tested on data from the same device. For example, the model trained on HydroCe achieves an EER of 12% on HydroCe test data but 27% on Geodisi test data. To further investigate why some hardware-specific models perform better on their own data, we summarized the training and testing data in Table 4, which reveals that the number of unique subjects in the training set has a stronger correlation with the EER. Therefore, weaker performance in some models could be related to the lower number of subjects available for learning features.
@@
General Model: We trained a general model using data from all devices. The results, presented in Table 2, show that the general model slightly outperforms hardware-specific models even on their own data. However, a follow-up question arises: does training a model on independent hardware-specific datasets contribute to better general models, or does the general model benefit from learning cross-device patterns in the current training dataset? To explore this, we created a smaller train dataset by retaining only the most frequent hardware per subject, ensuring that each subject exclusively has data recorded by a single device. The results in Table 3 indicate that the general model outperforms the Geodisi and BioSemi hardware-specific models but performs slightly worse than the HydroCe-specific model. Therefore, to some degree, we can conclude that combining smaller datasets with electrodes in the same positions can be beneficial for brainwave authentication, especially when large datasets are not available.
Cross Hardware: In real-world scenarios, individuals using brainwave-based authentication systems should be able to rely on various hardware devices. Therefore, subjects should still be authenticated when using different recording devices. Two key observations provide insight into this problem. First, the general model’s performance, when trained on a hardware-independent dataset, is presented in Table 3. It shows an error rate more than twice as high compared to Table 2, where the model was trained on a dataset including subjects with recordings from different hardwares. This suggests that providing cross-hardware samples during feature extractor training is essential for improving performance. Second, we filtered genuine scores based on unique enrollment and verification hardware pair combinations in the test dataset and calculated the EER for cross-hardware scenarios. The results show an EER of 13.25% for Geodisi-HydroCe pairs and 48.2% for BioSemi-HydroCe pairs and there were no HydroCe-BioSemi pairs in the test set. These results can be explained by the fact that only 7 subjects in the training set had cross BioSemi-HydroCe sessions, compared to 125 subjects who had cross Geodesic-HydroCe sessions. This could also explain the formation of separate clusters in the t-SNE visualization shown in Figure 5, where BioSemi forms a distinct cluster.
@@
Table 2. Table showing EER (%) with standard deviation ($\pm$ Std). The first column (D) lists the datasets, while the first row (M) lists the models. For example, the entry in the second row and third column (27.48 $\pm$ 10.39) represents the EER for the HydroCe model (train based on HydroCe data) evaluated on the Geodisi dataset. ”All” indicates results obtained using all data or models trained on data from all headsets.
@@
Geodisi $14.50\pm\text{\scriptsize 8.33}$ $27.48\pm\text{\scriptsize 10.39}$ $37.81\pm\text{\scriptsize 10.87}$ $13.37\pm\text{\scriptsize 7.13}$
HydroCe $30.34\pm\text{\scriptsize 11.15}$ $12.15\pm\text{\scriptsize 7.86}$ $34.63\pm\text{\scriptsize 10.37}$ $12.26\pm\text{\scriptsize 9.12}$
BioSemi $42.11\pm\text{\scriptsize 7.10}$ $39.74\pm\text{\scriptsize 9.22}$ $23.28\pm\text{\scriptsize 8.60}$ $20.10\pm\text{\scriptsize 9.62}$
All $27.93\pm\text{\scriptsize 12.16}$ $23.45\pm\text{\scriptsize 13.50}$ $30.71\pm\text{\scriptsize 12.42}$ $10.78\pm\text{\scriptsize 6.67}$
Table 3. Table showing EER (%) with standard deviation ($\pm$ Std). The first column (D) lists the datasets, and the first row (M) lists the models, where the training data for each subject is limited to their most frequently used hardware.
@@
Geodisi $21.93\pm\text{\scriptsize 10.40}$ $28.85\pm\text{\scriptsize 11.79}$ $34.43\pm\text{\scriptsize 7.46}$ $17.82\pm\text{\scriptsize 9.32}$
HydroCe $35.08\pm\text{\scriptsize 9.63}$ $12.49\pm\text{\scriptsize 9.23}$ $33.92\pm\text{\scriptsize 10.12}$ $13.82\pm\text{\scriptsize 9.60}$
BioSemi $43.99\pm\text{\scriptsize 5.94}$ $40.57\pm\text{\scriptsize 6.40}$ $25.52\pm\text{\scriptsize 8.15}$ $22.03\pm\text{\scriptsize 8.98}$
All $32.56\pm\text{\scriptsize 10.54}$ $26.42\pm\text{\scriptsize 16.29}$ $29.23\pm\text{\scriptsize 10.41}$ $24.37\pm\text{\scriptsize 17.22}$
Table 4. Summary of train datasets by hardware type. The table lists the number of sessions and subjects for each hardware type, with the corresponding percentages shown in parentheses.
@@
Table 5. Tradeoff between the number of samples (SN) and EER. Increasing the number of samples reduces the EER but increases the verification time, as shown in the table.
@@
Table 6. Tradeoff between the number of enrollment sessions (ES) and EER when using 1 verification sample (1V) and 4 verification samples (4V). The table also includes the number of available subjects (NAS) used in the evaluation.
@@
Table 7. EER (%) when using the same channels as consumer-grade devices. $E_{x}$ denotes the number of enrollment sessions, and $V_{x}$ denotes the number of verification sessions.
ES Baseline 4V E1 V1 E1 V4 E2 V1 E2 V4
Emotiv 13.97 10.74 10.44 6.01 8.79 4.48
DSI-VR300 16.55 12.70 13.49 8.56 11.94 6.87
Muse 2 21.94 17.74 18.07 12.42 17.58 11.19
4.3. Enhancing Performance
Samples from the same session are relatively close in the embedding space, but in some cases, they are more dispersed. For example, the pink triangle and green plus in Figure 6. Therefore utilizing multiple verification samples can improve results, albeit at the cost of additional verification time. Table 5 illustrates the tradeoff between the number of verification samples, EER, and verification time. The results show that using more samples reduces the error, but improvements beyond four samples show a slower pace in reducing error. Considering the time cost of approximately one second per sample, using four samples appears to be a reasonable balance for the averaging strategy.
Moreover, while comparing the enrollment session to all verification sessions provides valuable insights into brainwave authentication, real-world scenarios typically involve the first verification occurring immediately after enrollment. Additionally, verification sessions can potentially be used as supplementary data to enhance the enrollment process in the next verification attempt. Therefore, in Table 6, we consider the first session as enrollment and the second as verification and report the results. We observe an EER of 7.67% based on single verification sample and 4.23% for four verification samples, showing significant improvement compared to the previous 10.78% and 7.98%. Additionally, we investigate the effect of multiple enrollment sessions. The results suggest that having two enrollment sessions improves EER to 6.01% and 2.67%, but further enrollment sessions show diminishing returns. It is worth mentioning that the average time interval between two enrollment sessions was 28 days. Among the subjects, 88 had a time interval of 10 days or less, while 32 had session intervals of only one or two days.
4.4. Practicality with Consumer-Grade Channels

### v2

@@ IV-B1 Effect of the Number of Evaluation Subjects on EER
When comparing the results of different experiments using metrics such as EER, number of subjects in the evaluation is typically reported as metadata. A larger test set generates more impostor scores in zero-effort scenarios, making it valuable to investigate the relationship between the number of subjects in the evaluation set and the resulting error. This analysis provides insights for better comparisons of results across different studies or our follow-up experiments with smaller test sets, such as when evaluating performance using headset-specific recordings. To achieve this, we randomly select $N$ subjects, where $N$ ranges from 2 to 99 (starting at $N=2$ since at least one genuine user and one impostor are required). For each value of $N$, we compute the EER, repeat the procedure 50 times, and report the average EER per $N$.
@@
Figure 2 presents the results of the experiment. The analysis revealed no significant relationship between the number of subjects and the EER. These findings align with observations in face recognition systems. For example, Friedman et al. [51] investigated how test size affects identification and authentication tasks. They reported a linear decline in identification performance with the logarithm of enrollment size but found that the EER for authentication tasks remained stable. This stability occurs because the EER is based on the Receiver Operating Characteristic (ROC) curve, which balances the FAR and FRR derived from impostor and genuine score distributions. While larger subject pools better define these distributions, they do not significantly influence the ROC. Therefore, we can conclude that having a higher number of subjects in the test set does not necessarily make authentication more challenging. However, both scenarios indicate that test sets with a small number of participants exhibit higher deviations in EER, which can lead to unreliable or misleading results. This underscores the importance of using larger test sets to obtain more dependable evaluations.
Figure 4: t-SNE visualization of embeddings. Color represents the average cosine distance to the top 5 nearest samples from the same subject but different session. Warm colors indicate greater distances. Considering all samples, the average intra-class distance is 0.75 and the average inter-class distance, 0.97.
IV-B2 Correlation Between EER and Time Interval
@@
To investigate permanence, we first visualize the distribution of evaluation session intervals in Figure 3. The test data includes a total of 1,699 unique sessions available for 100 evaluation subjects, where the first session (or the oldest session) per subject is considered as the enrollment session (100 sessions). The results show that the highest concentration of sessions occurred within the first week, with 129 sessions (8%) recorded. By the end of the first month (weeks 0 to 3), this number increased to 560 sessions (35%), indicating that one-third of all sessions were conducted in the first month. The second month (weeks 4 to 7) accounted for 398 sessions (24.8%), while the third month (weeks 8 to 11) added 201 sessions (12.5%). Beyond the third month, session frequency declined further, with 182 sessions (11.3%) recorded from Month 4 to the end of the first year. In the second year, session numbers were 89 sessions (5.5%), and remained similar at 71 sessions (4.4%) in the third year. The remaining years combined accounted for 4.8%.
@@
To investigate the correlation between time intervals between sessions and authentication error, we extracted all possible pairs of sessions within specific time intervals and calculated the EER for each. Figure 3 shows the results, demonstrating a positive correlation between time interval and error rate, as expected. Verifying identity becomes harder over time. The EER after one day is 6.7%, increasing to 7.6% after seven days. However, within the second month, the EER rises to 7.2%, which remains lower than the previously reported average of 10.28%, indicating stable performance over a reasonable time interval. Beyond the second month, the EER gradually increases, reaching 14.3% after one year. Performance remains relatively stable in the second and third years. In the fourth and fifth years, performance deteriorates significantly, with EER values of 24.6% and 47.5%, respectively. However, due to the limited data for these time points, the results are not fully reliable. Nonetheless, even the doubling of EER after just one year suggests substantial degradation, indicating the need to update the enrollment set at shorter time intervals.
@@
Figure 5: t-SNE visualization of embeddings. Colors indicate different EEG headsets used for data collection.
IV-B3 Visualize Embedding Space
@@
To better understand how well users are represented in distinct clusters, we project the embeddings into a 2D space using t-Distributed Stochastic Neighbor Embedding (t-SNE) and visualize the results. Figure 4 shows that the metric learning approach successfully separates subjects: the average distance between a subject’s own samples is significantly lower (0.75) than the average distance to other subjects’ samples (0.97), indicating the model’s ability to capture individual differences. However, we observe two separate colonies within the embedding space, leading us to hypothesize that these colonies could be due to differences in the hardware used during data capture. This hypothesis is confirmed by Figure 5, which reveals that the data captured using BioSemi hardware is located separately from the rest, further illustrating the significant influence of hardware on the embedding distribution. The effect of hardware on the results will be discussed further in the next section.
@@
Figure 6: t-SNE visualization of embeddings for 20 randomly selected subjects, with each subject assigned a unique color for a clear distinction. Enrollment samples are depicted as circles without borders, while verification samples are represented by various shapes with black borders. Also, each unique shape corresponds to a distinct session for verification samples.
@@
Figure 7: Magnified view of the upper-right region of Figure 6, highlighting the light blue subject. Each session for the highlighted subject is assigned a unique color, while all other subjects are shown in gray for context.
@@
To investigate the embedding space while accounting for color and spatial limitations, we randomly selected 20 subjects to prevent overcrowding the embedding space with excessive data. [^7] BioSemi subjects were excluded to avoid large gaps in the plot and enhance detail visibility. Figure 6 illustrates the embedding space for these selected subjects, demonstrating that the feature extraction method effectively identifies subject identities. However, a session effect is evident, where samples from the same session tend to form distinct clusters. To better visualize this phenomenon, the upper-right region of Figure 6 is magnified in Figure 7, with separate colors indicating different sessions, further confirming the session-based clustering. The session effect on clustering may arise from environmental noise, changes in hair length, slight variations in the positioning of brainwave channels on the head, hardware differences, and natural changes in brainwave patterns over time due to factors such as aging, personal experiences, or mental states (e.g., stress or focus levels) throughout sessions.
IV-B4 Impact of Different Headsets on Performance
@@
Differences in EEG recording hardware and channel positions require further investigation to understand their impact on brainwave authentication. We set up to examine: (1) how well a model trained on one device performs on data from another, (2) whether models trained on multiple devices outperform device-specific models, and (3) how performs varies when enrollment and verification data come from different devices.
@@
Cross Model: To address the first question, we trained models using data from each specific headset and grouped the test data based on hardware to evaluate performance across models. The results, summarized in Table III, show that headset-specific models perform well only when tested on data from the same device. For example, the model trained on HydroCe achieves an EER of 9.2% on HydroCe test data but 16.1% on Geodesic test data. To further investigate why some hardware-specific models perform better on their own data, we summarized the training and testing data in Table V, which reveals that the number of unique subjects in the training set has a stronger correlation with the EER. Therefore, weaker performance in some models could be related to the lower number of subjects available for learning features.
@@
General Model: We trained a general model using data from all devices. The results in Table III show that the general model outperforms two hardware-specific models, even on their own data, but performs slightly worse than HydroCe. This comparison highlights two possible factors: the general model may benefit from learning cross-device patterns, or training on separate hardware-specific datasets may improve generalization. To examine this, we created a smaller training dataset by keeping only the most frequently used device per subject, ensuring each subject’s data comes from a single device. The results in Table IV show a similar pattern, although overall error rates are higher. Table V also shows that HydroCe has significantly more training samples than other devices. These findings suggest that combining smaller datasets with consistent electrode positions can be beneficial for brainwave authentication, especially when large datasets are not available.
Cross Device: In real-world scenarios, individuals using brainwave-based authentication systems may rely on different recording devices. Therefore, authentication systems should maintain performance when enrollment and verification data come from different hardware. To evaluate this, we filtered genuine scores based on unique enrollment-verification device pairs and calculated the EER for each cross-device setting. The results show an EER of 12.53% for Geodesic–HydroCe pairs and 12.56% for HydroCe–Geodesic pairs. In contrast, BioSemi–HydroCe and HydroCe–BioSemi pairs yield much higher EERs of 39.00% and 44.05%, respectively. For other combinations (e.g., Geodesic–BioSemi), no data were available in the evaluation set. The reason for this variation is that in the training set used for feature extraction, only 7 subjects had HydroCe–BioSemi session pairs and only 2 had BioSemi–HydroCe pairs out of 230 subjects, while 122 subjects had Geodesic–HydroCe session pairs. This limited overlap likely contributes to the poor generalization in the BioSemi–HydroCe condition and explains the formation of separate clusters in the t-SNE visualization shown in Figure 5, where BioSemi forms a distinct cluster.
@@
TABLE III: EER (%) and standard deviation ($\pm$ Std) for different datasets/models. The first column (D) lists the datasets, while the first row (M) lists the models. For example, the entry in the second row and third column (28.32 $\pm$ 11.7) represents the EER for the BioSemi model evaluated on the HydroCe dataset. ”All” indicates results obtained using all data or models trained on data from all headsets.
@@
[…]
TABLE IV: Table showing EER (%) with standard deviation ($\pm$ Std). The first column (D) lists the datasets, and the first row (M) lists the models, where the training data for each subject is limited to their most frequently used hardware.
@@
[…]
TABLE V: Summary of training datasets by hardware type, listing sessions and subjects with percentages in parentheses.
@@
IV-C Enhancing Performance
Samples from the same session are relatively close in the embedding space, but in some cases, they are more dispersed. For example, the pink session in Figure 6. Therefore utilizing multiple verification samples can improve results, albeit at the cost of additional verification time. Table VII illustrates the tradeoff between the number of verification samples, EER, and verification time. The results show that using more samples reduces the error, but improvements beyond four samples show a slower pace in reducing error. Considering the time cost of approximately one second per sample, using four samples appears to be a reasonable balance for the averaging strategy.
TABLE VI: Summary of state-of-the-art brain biometric verification solutions evaluated on multisession datasets
Publication Subjects Sessions Interval Verification Dataset Open-source Inter-subject EER (%)
[…]
Chaurasia et al. [14], 2024 54 108 (2) different days [^3] The paper did not specify the exact number of days, so we used the same terminology. However, it typically refers to a very short time interval. 1s 11.99
Proposed Method 345 6007 (17$\sim$ 17∼ 17) up to 66 months 1s 10.28
Moreover, while comparing the enrollment session to all verification sessions offers useful insights into brainwave authentication, real-world scenarios typically involve a single initial enrollment session followed by one verification attempt after some time. Upon successful login, it may be possible to update the enrollment set with new data from the verification attempt, which can then be used in future verifications. Therefore, in Table VIII, we consider the first session as enrollment and the second as verification and report the results. We observe an EER of 6.87% based on single verification sample and 3.04% for four verification samples, showing significant improvement compared to the previous 10.28% and 6.22%. Additionally, we investigate the effect of multiple enrollment sessions. The results suggest that having two enrollment sessions improves EER to 4.87% and 1.54%, but further enrollment sessions show diminishing returns. The average time interval between the two enrollment sessions was 28 days. Among the subjects, 88 had a session interval of 10 days or less, and 32 had intervals of only one or two days.
TABLE VII: Tradeoff between the number of samples (SN) and EER. Increasing the number of samples reduces the EER but increases the verification time, as shown in the table.
@@ IV-C Enhancing Performance
TABLE VIII: Tradeoff between the number of enrollment sessions (ES) and EER when using 1 verification sample (1V) and 4 verification samples (4V). The table also includes the number of available subjects (NAS) used in the evaluation.
@@
TABLE IX: EER (%) when using the same channels as consumer-grade devices. $E_{x}$ denotes the number of enrollment sessions, and $xV$ denotes the number of verification samples. For the first two columns (Baseline and 4V), all possible verification sessions were considered. However, for the remaining columns, only the first verification session immediately following the last enrollment session was used.
ES Baseline 4V E1 1V E1 4V E2 1V E2 4V
Emotiv 13.37 8.68 9.78 4.85 7.57 3.02
DSI-VR300 16.61 11.55 13.09 7.11 11.31 5.35
Muse 2 18.18 13.07 15.01 9.10 12.74 6.39
IV-D Practicality with Consumer-Grade Channels